A flywheel rotates with a constant retardation due to breaking from t=0to t=10 seconds. it made 400 revolutions . At time t=8 sec , it's angular velocity was 45πrad/sec. Determine

1 . Value of constant ratardation.

2. Total time taken to come to rest.

3. Total revolution made till it comes to rest.

Please let me know the answer as soon as possible . I need the answer today only please...SEND ME THE ANSWER FAST


1
Expert's answer
2021-05-05T07:46:06-0400

"\alpha =?\\
t_i= 0s\\
t_t = 10s\\
\omega= 0\ rad/s\\
\theta= 400\ rev= \dfrac{400}{2π}\ rad=\dfrac{200}{π}\ rad"



a.

"t= 8s\\
\omega = 45π\ rad/s"

from,

"\omega= \omega_o + \alpha t\\
0 =45π + \alpha (2)\\"

"\alpha = \dfrac{-45π}{2} = -22.5π\ rad/s²"



b.

from,

"\omega² = \omega_o²+ 2\alpha s\\
0 =\omega_o² + 2(-22.5π)(\dfrac{200}{π})\\

\omega_o= \sqrt{9000}= 94.87\ rad/s"


3.

Total revolution ("\theta") = "400\ rev= \dfrac{400}{2π}\ rad"

"= 63.7\ rad"